SOHO DAILY REPORT Operations Day – 105/2010 (15 April 10) Mission Day 5249 SOHO FLIGHT OPERATIONS TEAM, NASA/GSFC Ref. #:  SH-FOT-DR-100415 ==================================================================== PASS #1 Madrid    DSS65 0625-1020 PASS #2 Goldstone DSS27 1425-2255 ******************************************************************** PASS #1 Madrid    DSS65 0625-1020 Anomalies: None NRT: 0656-0959 Activities Summary: - SSR Dump - Collected MDI-M (0635-0652, 0854-1007) - OBT Distribution to CELIAS, SWAN - SSR Record ------------------ PASS #2 Goldstone DSS27 1425-2255 Anomalies: None NRT: 1502-2233 Activities Summary: - SSR Dump - Collected MDI-M (1435-1451, 1623-2242) - CRS Drift Estimation Update - Hx Limit Update - KKSEUCNT Reset - Cmd FSPAAD Latch Clear (Remains DETECTED) - PROS Branch B Pressure Check - CAE-A in the AOCS HW Config Word - SSR Record ------------------ -- End of Report --